# Cryptocurrency as an Investable Asset Class: Coming of Age

### v3

@@ 1 INTRODUCTION
Money has long been a central topic in economics—for example, see classic contributions of Tobin (1965) and Friedman (1968) to more recent advances. A large literature studied why money exists and how it functions—the role of money in economies with spatially separated agents (Townsend 1980), the coexistence of different types of money under different institutional restrictions (Wallace 1983), search-based models (Kiyotaki & Wright, 1989; Trejos & Wright, 1995; Shi, 1995; Lagos & Wright, 2005), mechanism-design approaches (Green & Zhou, 1998), the “money-as-memory” view (Kocherlakota, 1998), commodity and token money (Sargent & Wallace, 1983; Sargent, 2019), competing currencies (Kareken & Wallace, 1981), and the money-like safe assets (Gorton & Pennacchi 1990; Holmström & Tirole 1998).
@@
When practice advances faster than theory, it is natural to start asking empirical questions that can themselves later be used to shape theory. The paper of Liu & Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing. A sizable literature uses an asset pricing lens to study risk exposures, cross‑sectional return patterns, cash‑flow‑like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov & Schoar 2020; Liu, Tsyvinski & Wu 2022; Alvarez, Argente & Van Patten 2023; Sockin & Xiong 2023a; von Luckner, Reinhart & Rogoff 2023; Kogan et al. 2024).
@@
Many have viewed cryptocurrencies—and, not unjustifiably, many still view them—as hype, speculation, or fraud. When cryptocurrency first appeared it was easy to dismiss it: money engineered by an unknown author and operating outside of the traditional financial system. [^1] See, Foley, Karlsen & Putniņš (2019) and Griffin & Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations. It is worth noting that several important financial innovations, such as the limited‑liability company and mortgage‑backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
@@
Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi & Brunnermeier 2018; Leshno & Strack 2020; Biais et al. 2019; Kogan, Fanti & Viswanath 2022; Brunnermeier, James & Landau 2019; Benigno, Schilling & Uhlig 2022; Capponi, Jia & Wang 2025; Jermann & Xiang 2025), regulation (Cong & He 2019; Townsend 2020; Auer, Monnet & Shin 2025), digital asset pricing and valuation (Athey et al. 2016; Cong, Li & Wang 2021b; Biais et al. 2023a), transaction costs (Easley, O’Hara & Basu 2019; Cong, He & Li 2021a; Basu et al. 2023), central bank money (Auer et al. 2022; Whited, Wu & Xiao 2023; Schilling, Fernández-Villaverde & Uhlig 2024), securitization (Cong, Li & Wang 2022; Sockin & Xiong 2023a; Rogoff & You 2023; Goldstein, Gupta & Sverchkov 2024), safe assets (Ma, Zeng & Zhang 2025), international capital flow (von Luckner, Reinhart & Rogoff 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde & Sanches 2018) and market microstructure (Easley et al. 2024).
@@
We synthesize the rapidly growing empirical finance of cryptocurrencies into a set of seven stylized facts—empirical regularities that show how cryptocurrencies behave as assets, how their markets function, and what distinguishes them from traditional asset classes. These stylized facts help evaluate and revisit both classic and new theories of money, while remaining conditional on the available data and institutional environment to date and therefore potentially evolving as market structure, regulation, and adoption change. [^2] For complementary surveys on blockchain architecture, decentralized finance and central bank digital currencies (CBDC), see Kose et al. (2022); Makarov et al. (2022); Kose et al. (2023); Biais et al. (2023b); Adrian & Mancini-Griffoli (2021); Infante et al. (2023).
@@ 2 DATA
Second, for the construction of the crypto carry strategy of Schmeling, Schrimpf & Todorov (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025. The data on new addresses are from Liu, Tsyvinski & Wu (2021). Moreover, for the construction of the crypto arbitrage strategy of Borri & Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025. We convert these quotes into U.S. dollars using spot FX rates from Bloomberg. Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.333Finally, we obtain price indices for the MSCI global stock market, global high-yield and government bonds, a broad commodity index, gold, the U.S. dollar index, the Dow Jones US Real Estate Index, the U.S. constant maturity 10-year nominal and real bond yields from Bloomberg. The corresponding Bloomberg tickers are MXWO Index, LG30TRUU Index, LGTRTRUU Index, BCOM Index, XAU Curncy, DXY Curncy, DJUSRE Index, USGG10YR Index and GTII10 Govt.
@@
An important caveat concerns data integrity. A strand of the empirical cryptocurrency literature argues that parts of the historical crypto data—especially in early years—may reflect manipulation or unreliable reporting, including alleged price manipulation on Mt. Gox in 2013 (Gandal et al., 2018), inflated or “artificial” trading volume reported by some exchanges (Makarov & Schoar, 2020), and exchange-managed wash trading concentrated on “unregulated” exchanges and largely absent on “regulated” exchanges such as Bitstamp, Coinbase, and Gemini (Cong et al., 2023b), as well as trade reordering/front-running on decentralized exchanges (Daian et al., 2019). At the same time, on-chain evidence points to substantial cross-exchange flows: analyzing the Bitcoin blockchain, Makarov & Schoar (2021) show that a sizable share of volume to exchanges originates from other exchanges or large market participants, consistent with arbitrage activity that tends to limit persistent price deviations across major venues. While these concerns cannot be fully eliminated, we mitigate them in several ways. First, our CoinGecko sample starts after the alleged Mt. Gox manipulation episode in 2013. Second, our pre-/post-2020 split places emphasis on the most recent period, and the post-2020 subsample begins after the main evidence documented in Makarov & Schoar (2020) and Cong et al. (2023b). Third, the CoinGecko prices we use are volume-weighted averages across a curated set of large spot exchanges and, per CoinGecko’s documentation, [^5] See, https://www.coingecko.com/en/methodology. are constructed using algorithms designed to detect and ignore outliers; as of the end of our sample these include the following exchanges: Binance, Coinbase, Kraken, Crypto.com, Bitfinex, Bitstamp, Gemini, and Huobi. Finally, our market-capitalization screens and data filters are designed to downweight or exclude small, thinly traded coins where data quality issues are more likely to be severe.
@@ 3 THE SEVEN STYLIZED FACTS OF CRYPTOCURRENCY
Liu & Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable. Stylized fact 1 revisits and confirms this finding in a longer dataset through September 2025.777These performance measures are computed from price indices and thus reflect the return to the underlying asset, not the return to a particular investor after custody risk. Episodes such as exchange cybersecurity breaches or bankruptcies (e.g., Mt. Gox and FTX) can generate additional losses that are not reflected in spot price series.
@@
A complementary change concerns diversification. Table 1, Panel B, reports correlations between the coin market return and major asset classes. In the pre-2020 period, correlations with traditional assets are small and statistically insignificant, consistent with Liu & Tsyvinski (2021). In the post-2020 period, correlations rise sharply—most notably with global equities, increasing from 1% pre-2020 to 39% post-2020—and the coin market also becomes significantly correlated with commodities (27%), high-yield bonds (24%), government bonds (13%), and gold (18%). We also document a significant post-2020 correlation with changes in expected inflation (31%), measured as the first difference in the yield differential between nominal and inflation-indexed U.S. Treasury bonds with 10-year maturity. Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years.
The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong, Fang & Lin (2023) and Jiang et al. (2025). Several mechanisms are proposed to explain this phenomenon, including monetary policy, retail investor trading, and financialization. The literature has yet to reach a consensus on the cause of this important shift in the return dynamics of the cryptocurrency market. [^8] Results are similar using a broad U.S. stock market index such as the S&P 500 (2% pre-2020; 37% post-2020), while correlations with U.S. government bonds (rather than global government bonds) remain insignificant in the post-2020 period.
These correlation patterns matter directly for portfolio choice. In mean–variance settings, an asset’s contribution to diversification depends on its covariance with the rest of the portfolio rather than its standalone volatility (Markowitz, 1952). Consistent with the portfolio results in Liu & Tsyvinski (2021) and related Bayesian approaches (Duchin et al., 2025), a simple Black–Litterman calibration using post-2020 moments still implies non-trivial but modest allocations to the aggregate crypto market—about 3.1% for a representative retail portfolio and 5.5% for an institutional portfolio with a broader opportunity set. We emphasize that these magnitudes are back-of-the-envelope and depend on strong assumptions about expected returns, correlations, and investability (liquidity, custody, and regulatory constraints), and should therefore be interpreted as illustrative upper bounds on diversification benefits rather than investment advice.
@@ 3.2 Stylized fact 2: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value
A central question in asset pricing is whether systematic variation in expected returns can be explained by a parsimonious set of factors. In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama & French 1992)—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies. For cryptocurrencies, Liu, Tsyvinski & Wu (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns. Liu, Tsyvinski & Wu (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power. We revisit this evidence using our updated sample and show that a small number of predictors drive the cross-section of cryptocurrency returns over the longer sample and post-2020.
We construct smart beta strategies by sorting coins into quintile portfolios based on three established significant predictors: market capitalization (CSIZE), past two-week returns (CMOM), and the price-to-new-address ratio (CVALUE). [^9] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi, Babiak & Dickerson 2022), market segmentation (e.g., Borri & Shakhnov 2022), and higher-order moments (e.g., Borri et al. 2022b; Borri & Santucci de Magistris 2022). The return to each predictor is measured as the long-short spread between the highest and lowest quintiles (Portfolio 5 minus Portfolio 1).
Table 2, Panels A and B, report average excess returns to the quintile portfolios and the corresponding long-short strategies for the full sample and the post-2020 subsample. All three predictors generate statistically significant long-short returns in the full sample. The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu, Tsyvinski & Wu (2022). Cryptocurrency value is the weekly quintile portfolio of each bracket based on the price-to-new address ratio. High bracket is the low-value coins and low bracket is the high-value coins. The sorting is done in week $t$ and the returns are the value-weighted average in the following week.
@@
The C-4 structure of the crypto-market thus parallels the equity market, where a small number of equity-market factors—most notably size, value, momentum, and profitability—dominate the cross-section of expected returns. For example, Asness, Moskowitz & Pedersen (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities. The C-4 factor model of Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama & French (1992) and Carhart (1997) for equities. It is important to note that the predictors in cryptocurrency markets are only partly analogous to those in equity markets. CSIZE and CMOM closely resemble their equity counterparts, although crypto momentum is measured over a much shorter horizon (previous two weeks, versus the standard 12-to-2 months in equities). In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu, Tsyvinski & Wu (2021) and proxy CVALUE with the price-to-new-address ratio. [^10] Liu, Tsyvinski & Wu (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong, Li & Wang 2021b; Sockin & Xiong 2023a, b). More broadly, this stylized low-dimensional structure is consistent with the broader multifactor view of asset pricing developed for equities, but the relevant priced risks in cryptocurrency include distinct, crypto-specific components—including network/adoption-related fundamentals.
@@
A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”. Specifically, Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng, Giglio & Xiu 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model. In Table 3, we study long-short strategies based on these signals post-2020. We find that these strategies continue not to generate significant average returns in the recent sample. [^11] In a recent study, Fieberg et al. (2024) examine 20,736 research designs for 43 sorting variables in the cryptocurrency market. They find that the most prominent cryptocurrency factors, such as size and momentum, remain consistently robust across various specifications, but the others do not. These results show that the cross-section of cryptocurrency returns can be summarized by a small number of factors.

### v4

@@ 1 INTRODUCTION
Money has long been a central topic in economics—for example, see classic contributions of Tobin (1965) and Friedman (1968) to more recent advances. A large literature studied why money exists and how it functions—the role of money in economies with spatially separated agents (Townsend 1980), the coexistence of different types of money under different institutional restrictions (Wallace 1983), search-based models (Kiyotaki and Wright, 1989; Trejos and Wright, 1995; Shi, 1995; Lagos and Wright, 2005), mechanism-design approaches (Green and Zhou, 1998), the “money-as-memory” view (Kocherlakota, 1998), commodity and token money (Sargent and Wallace, 1983; Sargent, 2019), competing currencies (Kareken and Wallace, 1981), and the money-like safe assets (Gorton and Pennacchi 1990; Holmström and Tirole 1998).
@@
When practice advances faster than theory, it is natural to start asking empirical questions that can themselves later be used to shape theory. The paper of Liu and Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing. A sizable literature uses an asset pricing lens to study risk exposures, cross‑sectional return patterns, cash‑flow‑like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov and Schoar 2020; Liu, Tsyvinski, and Wu 2022; Alvarez, Argente, and Van Patten 2023; Sockin and Xiong 2023b; von Luckner, Reinhart, and Rogoff 2023; Kogan et al. 2024).
@@
Many have viewed cryptocurrencies—and, not unjustifiably, many still view them—as hype, speculation, or fraud. When cryptocurrency first appeared it was easy to dismiss it: money engineered by an unknown author and operating outside of the traditional financial system. [^1] See, Foley, Karlsen, and Putniņš (2019) and Griffin and Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations. It is worth noting that several important financial innovations, such as the limited‑liability company and mortgage‑backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
@@
Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi and Brunnermeier 2018; Leshno and Strack 2020; Biais et al. 2019; Kogan, Fanti, and Viswanath 2022; Brunnermeier, James, and Landau 2019; Benigno, Schilling, and Uhlig 2022; Capponi, Jia, and Wang 2025; Jermann and Xiang 2025), regulation (Cong and He 2019; Townsend 2020; Auer, Monnet, and Shin 2025), digital asset pricing and valuation (Athey et al. 2016; Cong, Li, and Wang 2021b; Biais et al. 2023), transaction costs (Easley, O’Hara, and Basu 2019; Cong, He, and Li 2021a; Basu et al. 2023), central bank money (Auer et al. 2022; Whited, Wu, and Xiao 2023; Schilling, Fernández-Villaverde, and Uhlig 2024), securitization (Cong, Li, and Wang 2022; Sockin and Xiong 2023b; Rogoff and You 2023; Goldstein, Gupta, and Sverchkov 2024), safe assets (Ma, Zeng, and Zhang 2025), international capital flow (von Luckner, Reinhart, and Rogoff 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde and Sanches 2018) and market microstructure (Easley et al. 2024).
@@
We synthesize the rapidly growing empirical finance of cryptocurrencies into a set of seven stylized facts—empirical regularities that show how cryptocurrencies behave as assets, how their markets function, and what distinguishes them from traditional asset classes. These stylized facts help evaluate and revisit both classic and new theories of money, while remaining conditional on the available data and institutional environment to date and therefore potentially evolving as market structure, regulation, and adoption change.
@@ 2 DATA
Second, for the construction of the crypto carry strategy of Schmeling, Schrimpf, and Todorov (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025. The data on new addresses are from Liu, Tsyvinski, and Wu (2021). Moreover, for the construction of the crypto arbitrage strategy of Borri and Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025. We convert these quotes into U.S. dollars using spot FX rates from Bloomberg. Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.222Finally, we obtain price indices for the MSCI global stock market, global high-yield and government bonds, a broad commodity index, gold, the U.S. dollar index, the Dow Jones US Real Estate Index, the U.S. constant maturity 10-year nominal and real bond yields from Bloomberg. The corresponding Bloomberg tickers are MXWO Index, LG30TRUU Index, LGTRTRUU Index, BCOM Index, XAU Curncy, DXY Curncy, DJUSRE Index, USGG10YR Index and GTII10 Govt.
@@
An important caveat concerns data integrity. A strand of the empirical cryptocurrency literature argues that parts of the historical crypto data—especially in early years—may reflect manipulation or unreliable reporting, including alleged price manipulation on Mt. Gox in 2013 (Gandal et al., 2018), inflated or “artificial” trading volume reported by some exchanges (Makarov and Schoar, 2020), and exchange-managed wash trading concentrated on “unregulated” exchanges and largely absent on “regulated” exchanges such as Bitstamp, Coinbase, and Gemini (Cong et al., 2023b), as well as trade reordering/front-running on decentralized exchanges (Daian et al., 2019). At the same time, on-chain evidence points to substantial cross-exchange flows: analyzing the Bitcoin blockchain, Makarov and Schoar (2021) show that a sizable share of volume to exchanges originates from other exchanges or large market participants, consistent with arbitrage activity that tends to limit persistent price deviations across major venues. While these concerns cannot be fully eliminated, we mitigate them in several ways. First, our CoinGecko sample starts after the alleged Mt. Gox manipulation episode in 2013. Second, our pre-/post-2020 split places emphasis on the most recent period, and the post-2020 subsample begins after the main evidence documented in Makarov and Schoar (2020) and Cong et al. (2023b). Third, the CoinGecko prices we use are volume-weighted averages across a curated set of large spot exchanges and, per CoinGecko’s documentation, [^4] See, https://www.coingecko.com/en/methodology. are constructed using algorithms designed to detect and ignore outliers; as of the end of our sample these include the following exchanges: Binance, Coinbase, Kraken, Crypto.com, Bitfinex, Bitstamp, Gemini, and Huobi. Finally, our market-capitalization screens and data filters are designed to downweight or exclude small, thinly traded coins where data quality issues are more likely to be severe.
@@ 3 THE SEVEN STYLIZED FACTS OF CRYPTOCURRENCY
Liu and Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable. Stylized fact 1 revisits and confirms this finding in a longer dataset through September 2025.666These performance measures are computed from price indices and thus reflect the return to the underlying asset, not the return to a particular investor after custody risk. Episodes such as exchange cybersecurity breaches or bankruptcies (e.g., Mt. Gox and FTX) can generate additional losses that are not reflected in spot price series.
@@
A complementary change concerns diversification. Table 1, Panel B, reports correlations between the coin market return and major asset classes. In the pre-2020 period, correlations with traditional assets are small and statistically insignificant, consistent with Liu and Tsyvinski (2021). In the post-2020 period, correlations rise sharply—most notably with global equities, increasing from 1% pre-2020 to 39% post-2020—and the coin market also becomes significantly correlated with commodities (27%), high-yield bonds (24%), government bonds (13%), and gold (18%). We also document a significant post-2020 correlation with changes in expected inflation (31%), measured as the first difference in the yield differential between nominal and inflation-indexed U.S. Treasury bonds with 10-year maturity. Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years.
The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong, Fang, and Lin (2023) and Jiang et al. (2025). Several mechanisms are proposed to explain this phenomenon, including monetary policy, retail investor trading, and financialization. The literature has yet to reach a consensus on the cause of this important shift in the return dynamics of the cryptocurrency market. [^7] Results are similar using a broad U.S. stock market index such as the S&P 500 (2% pre-2020; 37% post-2020), while correlations with U.S. government bonds (rather than global government bonds) remain insignificant in the post-2020 period.
These correlation patterns matter directly for portfolio choice. In mean–variance settings, an asset’s contribution to diversification depends on its covariance with the rest of the portfolio rather than its standalone volatility (Markowitz, 1952). Consistent with the portfolio results in Liu and Tsyvinski (2021) and related Bayesian approaches (Duchin et al., 2025), a simple Black–Litterman calibration using post-2020 moments still implies non-trivial but modest allocations to the aggregate crypto market—about 3.1% for a representative retail portfolio and 5.5% for an institutional portfolio with a broader opportunity set. We emphasize that these magnitudes are back-of-the-envelope and depend on strong assumptions about expected returns, correlations, and investability (liquidity, custody, and regulatory constraints), and should therefore be interpreted as illustrative upper bounds on diversification benefits rather than investment advice.
@@ 3.2 Stylized fact 2: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value
A central question in asset pricing is whether systematic variation in expected returns can be explained by a parsimonious set of factors. In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama and French 1992)—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies. For cryptocurrencies, Liu, Tsyvinski, and Wu (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns. Liu, Tsyvinski, and Wu (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power. We revisit this evidence using our updated sample and show that a small number of predictors drive the cross-section of cryptocurrency returns over the longer sample and post-2020.
We construct smart beta strategies by sorting coins into quintile portfolios based on three established significant predictors: market capitalization (CSIZE), past two-week returns (CMOM), and the price-to-new-address ratio (CVALUE). [^8] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi, Babiak, and Dickerson 2022), market segmentation (e.g., Borri and Shakhnov 2022), and higher-order moments (e.g., Borri et al. 2022b; Borri and Santucci de Magistris 2022). The return to each predictor is measured as the long-short spread between the highest and lowest quintiles (Portfolio 5 minus Portfolio 1).
Table 2, Panels A and B, report average excess returns to the quintile portfolios and the corresponding long-short strategies for the full sample and the post-2020 subsample. All three predictors generate statistically significant long-short returns in the full sample. The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu, Tsyvinski, and Wu (2022). Cryptocurrency value is the weekly quintile portfolio of each bracket based on the price-to-new address ratio. High bracket is the low-value coins and low bracket is the high-value coins. The sorting is done in week $t$ and the returns are the value-weighted average in the following week.
@@
The C-4 structure of the crypto-market thus parallels the equity market, where a small number of equity-market factors—most notably size, value, momentum, and profitability—dominate the cross-section of expected returns. For example, Asness, Moskowitz, and Pedersen (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities. The C-4 factor model of Liu, Tsyvinski, and Wu (2022) and Liu, Tsyvinski, and Wu (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama and French (1992) and Carhart (1997) for equities. It is important to note that the predictors in cryptocurrency markets are only partly analogous to those in equity markets. CSIZE and CMOM closely resemble their equity counterparts, although crypto momentum is measured over a much shorter horizon (previous two weeks, versus the standard 12-to-2 months in equities). In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu, Tsyvinski, and Wu (2021) and proxy CVALUE with the price-to-new-address ratio. [^9] Liu, Tsyvinski, and Wu (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong, Li, and Wang 2021b; Sockin and Xiong 2023b, a). More broadly, this stylized low-dimensional structure is consistent with the broader multifactor view of asset pricing developed for equities, but the relevant priced risks in cryptocurrency include distinct, crypto-specific components—including network/adoption-related fundamentals.
@@
A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”. Specifically, Liu, Tsyvinski, and Wu (2022) and Liu, Tsyvinski, and Wu (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng, Giglio, and Xiu 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model. In Table 3, we study long-short strategies based on these signals post-2020. We find that these strategies continue not to generate significant average returns in the recent sample. [^10] In a recent study, Fieberg et al. (2024) examine 20,736 research designs for 43 sorting variables in the cryptocurrency market. They find that the most prominent cryptocurrency factors, such as size and momentum, remain consistently robust across various specifications, but the others do not. These results show that the cross-section of cryptocurrency returns can be summarized by a small number of factors.